Show that if A and B are sets with the same cardinality, then |A|<=|B| and |B|<=|A|
1
Expert's answer
2020-10-18T17:06:59-0400

Given that , "A" and "B" are two set with same cardinality .i,e, "|A|=|B|"

Claim: "|A|\leq |B|" and "|B|\leq |A|"

Since, cardinality of "A" and "B" are equal .Therefore there exist a bijective function "f:A\rightarrow B"

Since "f" is one-one and "f(A)\sube B"

"\therefore |A| \leq |B|"

Again , "f^{-1}:B\rightarrow A" is a bijective map .

Therefore "f^{-1}" is one-one and "f^{-1}(B)\sube A"

Hence ,

"|B|\leq |A|"